consider a four-year project with the following information initial fixed asset investment = 580000; straight-line depreciation to zero over the four-year life; Zero salvage value; price = 42, variable costs= 29; fixed costs =255000, quantity sold =94000 units; tax rate = 24 percent

a. what is the degree of operating leverage at the given level of output?( Do not round intermediate calculations and round your answer to 4 decimal places)

b.What is the degree of operating leverage at the accounting break-even level of output? (Do not round Intermediate calculations and round your answer to 4 decimal places)

a.Degree of operating leverage

b. degree of operating leverage
Answer #1

Fixed assets = 580,000

Depreciction = 580,000 / 4 = 145000

Quantity sold = 94000

Price = 42

Sales = 42 * 94000 = 3948000

Variable cost = 29

Total variable cost = 29 * 94000 = 2726000

Fixed cost = 255000

EBITDA = 3948000 - 2726000 - 255000 = 967000

Depreciation = 145000

EBT =967000 - 145000 = 822000

Tax rate = 24%

Earnings after tax = 822000 * (1-0.24) = 624720

Breakeven point = Fixed cost / (price - variable cost)

Breakeven point = 255000 / (42 - 29) = 19615.38

DOL = quantity * ( price - variable cost ) / quantity * ( price - variable cost ) - Fixed cost

DOL = 19615.38 * 13 / ( 19615.38*13 ) - 255000

Since denominator is 0, At breakeven point, DOL is infinite

LET ME KNOW IF YOU HAVE ANY DOUBTS